Reports until 15:40, Thursday 19 October 2023
Lasers Amplifiers
Torrey Cullen - posted 15:40, Thursday 19 October 2023 - last comment - 17:13, Friday 20 October 2023(11336)
Input coupler test

[Alex Torrey]

We did a quick test of the input couplers we have on hand. Initially couldn't get more than 1% through the fiber. Plopped a f=1m lens about a meter away from the input coupler and now get ~80%. Will com eup with a proper MM solution for this tomorrow. 

 

Images attached to this report
Comments related to this report
Torrey Cullen - 17:13, Friday 20 October 2023 (11338)

The input coupler was not in its final position so we moved it and recovered ~80% transmission through the fiber. We discovered that how the input couplers are mounted currently is not fully recoverable. We will try turning on the SHG on monday and when the proper mounts adapters and mounts arrive (tuesday) we will put the input couplers in the new mounts and recover the fiber alignment.